# Epistemic Boundaries and Quantum Indeterminacy: What Local Observers Can (Not) Predict

### v2

@@ Results
Adlam and Rovelli follow a similar line of reasoning. In their recent work the authors define a notion of ‘cross-perspective links’ which is supposed to establish a consistency of measurement outcomes among different observers in relational quantum mechanics. [2]). That is, when some observer obtains the value of a physical variable producing a record of that value, and another observer subsequently measures that same record, the two results should match.
This goes in the direction of reliable intersubjectivity as defined in this work. Note, however, that Adlam and Rovelli’s definition is ambiguous in terms of whether the second measurement result is a reliable record of the first measurement result. The two values can match after the second interaction, but this doesn’t imply that at the time of the first measurement the outcome had the same value. Moreover, the claim established in this paper might provide a counterargument to ‘cross-perspective links’. For their theory seems to accept all the other assumptions of our result which in turn implies that reliable intersubjectivity cannot hold in relational quantum mechanics.
The assumptions mentioned apply for a single choice of measurement only (i.e. when Wigner checks the friend’s reading), and therefore prove insufficient to account for the more general assumption $(RI)$. Alice performs measurements in an arbitrary basis, for which the same question arises: What can Maggie predict about Alice’s outcomes conditioned on $z$?
$(RI)$ and $(PA)$ stay away from any commitment to whether Maggie’s outcome $z$ has anything to do with the ‘actual’ state of the physical system when it’s a superposed state according to Alice. Upon $(RI)$ the variable $z$ can be objectively communicated among different observers, but it is only by virtue of $(PA)$ that $z$ assumes the meaning of a variable ‘completing’ the quantum state. Moreover, we suppose $z$ to signify an arbitrary predictive advantage. That is, $z$ is not constraint to deterministic predictions for one particular setting (e.g. the ‘record checking’ measurement), or in fact to any deterministic prediction at all.
Assumption 4 (Signal-Locality (SL)).
The outcomes of an observer $A$ are conditionally independent of the settings of another space-like separated observer $B$ even when conditioned on other outcomes not in the future light-cone of $A$.
Upon reliable intersubjectivity the record of variable $z$ can be observed by both Alice and Maggie. Hence, both observers will objectively agree on the single value of $z$, and the process of Alice attaining this record has no effect on $z$. As a result, since Alice is in a position to access and employ $z$, the signal-locality condition has to hold when conditioned on the variable $z$. For otherwise, Alice could choose her settings depending on $z$ to send superluminal signals to Bob.
@@
and this is non-trivial in the case where $z$ provides predictive advantage via Assumption 2. Signal-locality is a weaker notion of locality than Bell’s parameter independence for hidden variables since if one is to propose the existence of such variables they may still be parameter dependent without $z$ leading to violation of signal-locality. Secondly, the assumption applies to actually observed events of Maggie. All the variables in the statement of the assumption are manifest configurations in the supervenience domain. Signal-locality provides the last assumption necessary for the main result, together with one auxiliary condition, namely
@@
Measurement settings are independent of outcomes outside their future light-cone.
In the context of our setup this reads,
$p(z|a,b)=p(z),$ (2.5)
also known as the ‘freedom of choice’, or ‘no-superdeterminism’ condition, the assumption ensures that Alice and Bob can choose their measurement settings independently of the magic prediction box.
We are now ready to state the main result of this work which is the claim that assumptions (1)–(5) cannot all jointly be true. That is,
$\bm{(SD)\wedge(PA)\wedge(RI)\wedge(SL)\wedge(NC)\rightarrow\bot.}$ (2.6)
The proof of this claim is found in §3.
Discussion
This work contributes to the empirical completeness problem of quantum theory, i.e. the question whether theories could exist with predictive advantage while conforming to the Born rule on average. Drawing on quantum non-locality by generalised Bell inequalities for maximally entangled states, we have shown that whenever there would exist predictive advantage, observers could superluminally signal.
But the result goes beyond the mere claim that quantum mechanics is predictively complete. It establishes under what precise conditions quantum probabilities impose epistemic boundaries between observers, notably what we introduced as reliably intersubjectivity. The result highlights that this is a necessary assumption to be made in order to arrive at empirical completeness.
Moreover, this work advances existing related results. For example, it addresses some modern results involving Wigner’s friend type scenarios. As mentioned, the assumption of $(AOE)$ in the no-go theorem by Bong et al., can be reframed as a special case of reliable intersubjectivity and predictive advantage. Their claim follows from restricting the definition of predictive advantage to the case where Maggie’s probabilities conditioned on $z$ yield a deterministic prediction for one of three measurement settings employed in their setup. That is, $(AOE)$ amounts to reliable intersubjectivity together with a special version of predictive advantage. In the context of this work Bong et al.’s no-go theorem thus shows that a certain measurement scenario can be found ruling out deterministic predictive advantage for one of three measurement settings (barring the possibility of violating any of the other assumptions in the theorem).
@@ Discussion
In a very recent paper by Moreno et al. the authors tried to quantify $(AOE)$ and showed that ‘observed facts are maximally non-absolute’ by invoking versions of generalised Bell inequalities.[34] This is in the spirit of our claim, but the conclusion of our result is about the predictability of observers rather than the ‘absoluteness of facts’: If reliable intersubjectivity—i.e. the more general version of $(AOE)$—doesn’t hold, then Maggie can in principle have predictive advantage at the cost of it being ‘private’ and thus ‘non-absolute’. This kind of ‘improved predictivity’ therefore introduces an epistemic boundary between Maggie and the other observers she makes predictions about. [^9] Note further that Moreno et al introduce a relaxation of $(AOE)$ which they call $\epsilon$-AOE and it is the assumption that for the special measurement that Cavalcanti consider (the record checking basis) the probability that the two outcomes are equal is not 1 but 1-epsilon. This again can be related to a special choice of predictive advantage where for the record checking basis the predictions are probabilistic. Thus, their assumption is related to predictive advantage but not as general as how it is defined here.
@@
The proof has been analysed and criticised, e.g. in [24, 29, 28, 26, 20]. The criticism is mainly about Colbeck and Renner’s assumptions implying parameter independence (which is much stronger than assuming signal-locality alone), the status of ‘free will’ and its relationship to the central premise, and, moreover, lacking the foundation for addressing individual systems for which much stronger assumptions seem necessary.
@@
Furthermore, Colbeck and Renner do not address the epistemic nature of the variables (here Maggie’s records $z$), and, importantly, how they relate to the variables of other observers—as is established here by explicitly assuming $(RI)$. This confuses their analysis, for it is hence not entirely clear if the result pertains to controllable or uncontrollable variables. The claims crucially hinge on whether there exists an epistemic boundary between Maggie and Alice. [^12] Consequently, it depends on the details of what is assumed about the variables’ relationship to one another whether the locality assumption amounts to signal-locality or parameter-independence, and whether the variables are even uniquely defined for all observers. In the accounts of Hermens and Leifer, the theorem is restated in the context of ontological models, providing an unambiguous underpinning of the result. There the ‘free-choice’ assumption (FR) straightforwardly translates to parameter independence (and settings independence). This is implied by the fact that the relevant variables enter the story as the full ontic (non-manifest) state $z=\lambda$ of a preparation determining all outcome probabilities of arbitrary measurements. This seems to be the view as standardly conceived (see, for instance, [29, 20, 24, 28, 26]). Then again, the problem remains whether $z$ has anything to do with an actual measurement record that can be communicated to other observers.
@@
For example, the authors purport that their assumptions rule out pilot wave theory—which is true since (FR) does imply parameter independence (and is violated in Bohmian theories). But it is worth mentioning that pilot wave theory is signal-local while still being parameter dependent. In fact, hidden variable theories are not a priori ruled out by the assumptions made in our approach. Regarding pilot wave theory our claim simply shows that the Bohmian configurations (if assumed to exist) cannot be jointly known by Maggie and Alice (although $z$ may still represent a Bohmian particle configuration) if they are to deterministically fix all measurement outcomes. [^13] See also the idea of a coarse-grained hidden variable in Ghirardi and Romano’s approach, which resembles the thought that the $z$ variables don’t necessarily refer to the full ontic states [21].
@@
Intriguingly, Maggie can in principle have knowledge of the outcomes of observables corresponding to non-commuting operators, but wouldn’t be able to communicate that to Alice. As long as reliable intersubjectivity is violated, she can predict deterministically the outcomes $x$ of Alice’s measurement for arbitrary settings $a$. In a hidden variable setting (e.g. see below in pilot wave theory) this is in fact possible whenever before the observer’s memory is erased when interacting with the system. The result shows that Maggie cannot communicate that knowledge to Alice since that would disturb Maggie’s record $z$ in order to reproduce the quantum statistics. In Everettian quantum mechanics a similar story can be told. However, there is no disturbance of the memory involved since the possible outcomes exist already in different branches, and Maggie’s wave function branch will match the corresponding branch of Alice. In this case, reliable intersubjectivity is violated by introducing non-unique outcomes, i.e. $z$ doesn’t posses a single definite value.
Our result also connects to the quantum measurement problem. It is thus interesting to look at some of the quantum interpretations in the context of reliable intersubjectivity and predictive advantage.
@@
For example, in (equilibrium) pilot wave theory [6, 22] the particle configurations are non-quantum variables in the sense that in general they do not satisfy the quantum decorrelation principle—e.g. the particle having a definite position doesn’t imply that the corresponding wavefunction has very narrowly peaked support. As it turns out, it is no contradiction for an observer to obtain Bohmian particle configurations, even in the cases where they behave non-classically. That is, in a Bohmian world it is in principle possible for Maggie to possess a magic prediction box that spits out the Bohmian initial configurations of the setup, therefore having deterministic predictive advantage. But upon signal-locality—which is satisfied in pilot wave theory on the level of manipulable variables—her records cannot be reliably reported to Alice. This follows from the observation that A) Maggie can in principle measure the Bohmian configurations represented by $z$, and they would be perfectly correlated to Alice’s measurement outcomes, but B) if Maggie’s records could be reliably communicated to Alice, then Alice could use them to superluminally signal to Bob since Bohmian configurations are non-local. Thus, although Maggie may be able to observe them initially and make an improved prediction, her records must be disturbed through interaction with Alice in order to recover the correct quantum statistics on the level of Alice. Interestingly, our result implies that this holds for any kind of super-quantum predictions. In the context of hidden variable theories the interaction between Maggie and Alice thus always introduces disturbance whenever the variable $z$ represents any degree of predictive advantage. Reliable intersubjectivity must thus be false in such theories to restore signal-locality on the operational level. Brukner et al. came to a similar conclusion for a special measurement scenario involving two measurements, and showed a no-go result for ‘the persistent reality of Wigner’s friend’s memory’.[3]
@@
Whenever Maggie’s variable would constitute some pre-determination of Alice’s outcomes beyond the quantum probabilities, its value will be concealed from Alice. They cannot live in a mutual ‘reality’ where the record $z$ exists for both simultaneously. This epistemic boundary between observers is the mechanism by which ‘hidden variables’ are hidden.
@@
In Everettian quantum mechanics [17, 25], reliable intersubjectivity is violated due to measurement records not being unique. Again the question arises if this could constitute predictive advantage. The variables that are in principle available to an observer like Maggie are the total wave function of the setup and knowledge of the branch she is in. Conditioned on the total wave function alone Maggie’s means to predict Alice’s outcomes are limited by the Born probabilities. Thus, the question is whether knowledge of the subjective branch could improve her predictions. In certain scenarios Maggie can indeed have relative predictive advantage. For instance, Maggie may have measured a quantum system with two outcomes where after the interaction there exist two copies of Maggie having recorded one of the outcomes. The total wavefunction is still a superposition of those two branches, but if Alice measures in the same basis—i.e. checking Maggie’s record—each Maggie can predict with certainty that there will be an Alice observing the same outcome, and, thus each Maggie has relative predictive advantage. It would be interesting to see what the conclusions are for other choices of Alice’s basis. We leave that for further investigations. We suspect that analogous conclusions can be drawn in Rovelli’s ‘relational quantum mechanics’ [37] and pragmatist views of quantum mechanics like QBism [23].
@@
The epistemic boundaries between observers resulting from a failure of $(RI)$ afford an analogy in the spirit of Laplace’s demon in classical physics. The demon has to be an isolated creature external to the universe that it is describing in order to avoid logical inconsistencies whenever it is itself part of the system. Although there exists no direct link to a self-referential problem in the quantum case, the situation is similar in that quantum observers could exist with predictive advantage (even deterministic) but which are epistemically separated from the system and other observers. Moreover, when the different observers interact, in order to keep the epistemic boundary intact, reliable intersubjectivity is violated such that either observed records can have well-defined values but Maggie’s variable $z$ (and her memory states) have to flip, or, the observed records of $z$ are subjective and thus don’t assume well-defined values, but the variable $z$ is preserved when communicating with Alice.
Methods

### v3

@@ 3.1 Formalisation of the Assumptions
Adlam and Rovelli follow a similar line of reasoning. In their recent work, the authors define a notion of ‘cross-perspective links’ which is supposed to establish a consistency of measurement outcomes among different observers in relational quantum mechanics. [26]). That is, when some observer obtains the value of a physical variable producing a record of that value, and another observer subsequently measures that same record, the two results should match.
This goes in the direction of reliable intersubjectivity as defined in this work. Note, however, that Adlam and Rovelli’s definition is ambiguous in terms of whether the second measurement result is a reliable record of the first measurement result. The two values can match after the second interaction, but this doesn’t imply that at the time of the first measurement, the outcome had the same value. Moreover, the claim established in this paper might provide a counterargument to ‘cross-perspective links’. For their theory seems to accept all the other assumptions of our result which in turn implies that reliable intersubjectivity cannot hold in relational quantum mechanics.
The assumptions mentioned apply for a single choice of measurement only (i.e. when Wigner checks the friend’s reading) and therefore prove insufficient to account for the more general assumption $(RI)$. Alice performs measurements in an arbitrary basis, for which the same question arises: What can Maggie predict about Alice’s outcomes conditioned on $z$?
$(RI)$ and $(PA)$ stay away from any commitment to whether Maggie’s outcome $z$ has anything to do with the ‘actual’ state of the physical system when it’s a superposed state according to Alice. Upon $(RI)$ the variable $z$ can be objectively communicated among different observers, but it is only by virtue of $(PA)$ that $z$ assumes the meaning of a variable ‘completing’ the quantum state. Furthermore, we suppose $z$ to signify an arbitrary predictive advantage. That is, $z$ is not constrained to deterministic predictions for one particular setting (e.g. the ‘record checking’ measurement), or in fact to any deterministic prediction at all.
Assumption 4 (No-Signalling (SL)).
The outcomes of an observer $A$ are conditionally independent of the settings of another space-like separated observer $B$ even when conditioned on other outcomes not in the future light cone of $A$.
Upon reliable intersubjectivity, the record of variable $z$ can be observed by both Alice and Maggie. Hence, both observers will objectively agree on the single value of $z$, and the process of Alice attaining this record does not affect $z$. As a result, since Alice is in a position to access and employ $z$, the no-signalling condition has to hold when conditioned on the variable $z$. Otherwise, Alice could choose her settings depending on $z$ to send superluminal signals to Bob.
@@
and this is non-trivial in the case where $z$ provides predictive advantage via Assumption 2. No-signalling is a weaker notion of locality than Bell’s parameter independence for hidden variables since if one is to propose the existence of such variables they may still be parameter dependent without $z$ leading to violation of no-signalling. Secondly, the assumption applies to actually observed events of Maggie. All the variables in the statement of the assumption are manifest configurations in the supervenience domain. No-signalling provides the last assumption necessary for the main result, together with one auxiliary condition, namely
@@
Measurement settings are independent of outcomes outside their future light cone.
Alice and Bob should be able to choose their measurements freely and independently of the quantum system (which includes the specification of Maggie’s variable $z$). It can be assumed that the settings $a$ and $b$ are outside of the future light cone of $z$. That is, we have $p(a,b|z)=p(a,b)$, and by the definition of conditional probabilities the assumption $(NC)$ reads
$p(z|a,b)=p(z).$ (7)
In the context of hidden variables theories $(NC)$ is also known as the ‘freedom of choice’, ‘no-superdeterminism’ condition, or setting independence. [^12] For further considerations on this assumption see [27].
We are now ready to state the main result of this work, which is the claim that assumptions (1)–(5) cannot all jointly be true. That is,
$\bm{(SD)\wedge(PA)\wedge(RI)\wedge(SL)\wedge(NC)\rightarrow\bot.}$ (8)
The proof of this claim is found in Appendix A.
4 Discussion
Drawing on quantum non-locality by generalised Bell inequalities for maximally entangled states, we have shown that whenever there would exist any predictive advantage, observers could superluminally signal unless reliable intersubjectivity (or no-conspiracy) is violated.
Here we shall discuss related questions and results. We first draw conclusions for the ‘absoluteness of observed events’. We then comment on the Colbeck-Renner theorem and then discuss the possibility of violating reliable intersubjectivity. We conclude with a brief discussion of pilot wave theory, Everettian quantum mechanics, and a note on Laplace’s demon.
Absoluteness of Observed Events.
This work advances existing related results involving Wigner’s friend-type scenarios. As mentioned, the assumption of $(AOE)$ in the no-go theorem by Cavalcanti et al. [8] can be reframed as a special case of reliable intersubjectivity and predictive advantage. Their claim follows from restricting the definition of predictive advantage to the case where Maggie’s probabilities conditioned on $z$ yield a deterministic prediction for one of three measurement settings employed in their setup. That is, $(AOE)$ amounts to reliable intersubjectivity together with a version of predictive advantage. In the context of this work, Bong et al.’s no-go theorem shows that a certain measurement scenario can be found ruling out deterministic predictive advantage for one of three measurement settings (barring the possibility of violating any of the other assumptions in the theorem).
@@ 4 Discussion
In a very recent paper by Moreno et al., the authors tried to quantify $(AOE)$ and showed that ‘observed facts are maximally non-absolute’ by invoking versions of generalised Bell inequalities.[28] This is in the spirit of our claim, but the conclusion of our result is about the predictability of observers rather than the ‘absoluteness of facts’: If reliable intersubjectivity—i.e. the more general version of $(AOE)$—doesn’t hold, then Maggie can in principle have predictive advantage at the cost of it being private and hence ‘non-absolute’. This kind of improved predictivity therefore introduces an epistemic boundary between Maggie and other observers. [^13] Note further that Moreno et al introduce a relaxation of $(AOE)$ which they call $\epsilon$-AOE and it is the assumption that for the special measurement that Cavalcanti consider (the record checking basis) the probability that the two outcomes are equal is not 1 but 1-epsilon. This again can be related to a special choice of predictive advantage where for the record checking basis the predictions are probabilistic. Thus, their assumption is related to predictive advantage but not as general as how it is defined here.
Colbeck-Renner Theorem.
@@
The proof has been analysed and criticised, e.g. in [30, 31, 32, 33, 34]. The criticism is mainly about Colbeck and Renner’s assumptions implying parameter independence (which is much stronger than assuming no-signalling alone), the status of ‘free will’ and its relationship to the central premise, and, moreover, lacking the foundation for addressing individual systems for which much stronger assumptions seem necessary.
@@
Furthermore, Colbeck and Renner do not address the epistemic nature of the variables (here Maggie’s records $z$), and, importantly, how they relate to the variables of other observers—as is established here by explicitly assuming $(RI)$. This confuses their analysis, for it is hence not entirely clear if the result pertains to controllable or uncontrollable variables. The claims crucially hinge on whether there exists an epistemic boundary between Maggie and Alice. [^16] Consequently, it depends on the details of what is assumed about the variables’ relationship to one another whether the locality assumption amounts to no-signalling or parameter-independence, and whether the variables are even uniquely defined for all observers. In the accounts of Hermens and Leifer, the theorem is restated in the context of ontological models, providing an unambiguous underpinning of the result. There the ‘free-choice’ assumption (FR) straightforwardly translates to parameter independence (and settings independence). This is implied by the fact that the relevant variables enter the story as the full ontic (non-manifest) state $z=\lambda$ of a preparation determining all outcome probabilities of arbitrary measurements. This seems to be the view as standardly conceived (see, for instance, [31, 34, 30, 32, 33]). Then again, the problem remains whether $z$ has anything to do with an actual measurement record that can be communicated to other observers.
@@
For example, the authors purport that their assumptions rule out pilot wave theory—which is true since (FR) does imply parameter independence (and is violated in Bohmian theories). But it is worth mentioning that pilot wave theory is no-signalling while still being parameter dependent. Hidden variable theories are not a priori ruled out by the assumptions made in our approach. Regarding pilot wave theory our claim shows that the Bohmian configurations (if assumed to exist) cannot be jointly known by Maggie and Alice (although $z$ may still represent a Bohmian particle configuration) if they are to deterministically fix all measurement outcomes. [^17] See also the idea of a coarse-grained hidden variable in Ghirardi and Romano’s approach, which resembles the thought that the $z$ variables don’t necessarily refer to the full ontic states [35].
Violations of Reliable Intersubjectivity.
@@
Intriguingly, Maggie can in principle know the outcomes of observables corresponding to non-commuting operators, but wouldn’t be able to communicate that to Alice. As long as reliable intersubjectivity is violated, she can predict deterministically the outcomes $x$ of Alice’s measurement for arbitrary settings $a$ (see pilot wave theory below). Nevertheless, the result shows that Maggie cannot communicate that knowledge to Alice since that would invariably disturb Maggie’s record of $z$. In Everettian quantum mechanics, a similar story can be told (see Everettian Quantum Mechanics below). However, there is no disturbance of the memory involved since the possible outcomes exist already in different branches, and Maggie’s wave function branch will match the corresponding branch of Alice. In this case, reliable intersubjectivity is violated by introducing non-unique outcomes, i.e. $z$ doesn’t possess a single definite value.
Pilot Wave Theory.
In (equilibrium) pilot wave theory [36, 37] the particle configurations are non-quantum variables in the sense that in general they do not satisfy the quantum decorrelation principle—e.g. the particle having a definite position doesn’t imply that the corresponding wavefunction has very narrowly peaked support.
@@
As it turns out, it is no contradiction for an observer to obtain Bohmian particle configurations, even in the cases where they behave non-classically. That is, in a Bohmian world it is in principle possible for Maggie to possess a magic prediction box that spits out the Bohmian initial configurations of the setup, therefore having deterministic predictive advantage. But upon no-signalling—which is satisfied in pilot wave theory on the level of manipulable variables—her records cannot be reliably reported to Alice. This follows from the observation that A) Maggie can in principle measure the Bohmian configurations represented by $z$, and they would be perfectly correlated to Alice’s measurement outcomes, but B) if Maggie’s records could be reliably communicated to Alice, then Alice could use them to signal superluminally to Bob (since Bohmian configurations are non-local). Thus, although Maggie may be able to observe them initially and make an improved prediction, her records must be disturbed through interaction with Alice to recover the correct quantum statistics on the level of Alice. This should hold for any kind of super-quantum predictions. In the context of hidden variable theories the interaction between Maggie and Alice thus always introduces disturbance whenever the variable $z$ represents any degree of predictive advantage. Reliable intersubjectivity must thus be false in such theories to restore no-signalling on the operational level. Brukner et al. came to a similar conclusion for a special measurement scenario involving two measurements, and showed a no-go result for ‘the persistent reality of Wigner’s friend’s memory’.[11]
@@
Whenever Maggie’s variable would constitute some pre-determination of Alice’s outcomes beyond the quantum probabilities, its value will be concealed from Alice. They cannot both live in a mutual ‘reality’ where the record $z$ exists for both simultaneously. This epistemic boundary between observers is the mechanism by which ‘hidden variables’ are hidden.
Everettian Quantum Mechanics.
@@
In Everettian quantum mechanics [38, 39], reliable intersubjectivity is violated due to measurement records not being unique. Again the question arises if this could constitute predictive advantage. The variables that are in principle available to an observer like Maggie are the total wave function of the setup and knowledge of the branch she is in. Conditioned on the total wave function alone Maggie’s means to predict Alice’s outcomes are limited by the Born probabilities. Thus, the question is whether knowledge of the subjective branch could improve her predictions. In certain scenarios Maggie can indeed have relative predictive advantage. For instance, Maggie may have measured a quantum system with two outcomes where after the interaction there exist two copies of Maggie having recorded one of the outcomes. The total wavefunction is still a superposition of those two branches, but if Alice measures in the same basis—i.e. checking Maggie’s record—each Maggie can predict with certainty that there will be an Alice observing the same outcome, and, thus each Maggie has relative predictive advantage. It would be interesting to see what the conclusions are for other choices of Alice’s basis. We leave that for further investigations.
We suspect that analogous conclusions can be drawn in Rovelli’s ‘relational quantum mechanics’ [40] and pragmatist views of quantum mechanics like QBism [41].
Laplace’s Demon.
@@
The epistemic boundaries between observers resulting from a failure of $(RI)$ afford an analogy in the spirit of Laplace’s demon in classical physics. The demon has to be an isolated creature external to the universe that it is describing in order to avoid logical inconsistencies whenever it is itself part of the system. Although there exists no direct link to a self-referential problem in the quantum case, the situation is similar in that quantum observers could exist with predictive advantage (even deterministic) but who are epistemically separated from the system and other observers. Moreover, when the different observers interact, to keep the epistemic boundary intact, reliable intersubjectivity is violated such that either observed records can have well-defined values and Maggie’s variable $z$ (and her memory states) has to flip, or, the observed records of $z$ are subjective and not well-defined values, but the variable $z$ is preserved when communicated to Alice.
5 Conclusions
Can post-quantum theories have predictive advantage whilst conforming to the Born rule on average? We showed that unless epistemic boundaries exist between observers, the answer is no: quantum uncertainty is an unavoidable limit to Nature’s predictability. For any hypothetical predictive advantage, there exists a quantum measurement scenario with a finite number of settings which implies that predictive advantage would violate no-signalling. Giving up the assumption of ‘reliable intersubjectivity’ does in principle enable (subjective) predictive advantage (and even knowledge about the outcomes of non-commuting observables), but implies an epistemic boundary between different quantum observers.
To our knowledge, this is the first rigorous result on predictive completeness addressing locality and the mutual epistemic reality of macroscopic observers.
We left open the question whether (operational) principles—other than no-signalling—can be invoked for a claim about empirical completeness. For example, these could be dynamical or methodological assumptions about the prediction improving variables. Moreover, no general proof is currently available for arbitrary single systems. The impossibility of predictive advantage for single systems may be established on grounds of non-contextuality—as a generalisation of locality.
In sum, if we assume that observers can reliably communicate and can’t send superluminal signals, the contemporary mainstream view that quantum predictions are irreducibly probabilistic becomes a precise theorem. Consequently, not only can quantum mechanical measurement outcomes not be predicted with certainty, the no-signalling predictions of any theory must be exactly the quantum probabilities.
Acknowledgments
Appendix A Methods